Kevin invested his savings in two investment funds. The amount he invested in Fund A was $3000 less than the amount he invested in Fund B. Fund A returned a
8% profit and Fund B returned a 5% profit. How much did he invest in Fund A, if the total profit from the two funds together was $1320?

Answer by lwsshak3(11628)

Kevin invested his savings in two investment funds. The amount he invested in Fund A was $3000 less than the amount he invested in Fund B. Fund A returned a
8% profit and Fund B returned a 5% profit. How much did he invest in Fund A, if the total profit from the two funds together was $1320?
***
let x= amount invested in Fund B
x-3000= amount invested in Fund A
..
8%((x-3000)+5%x=1320
.08x-240+.05x=1320
.13x=1560
x=12000
x-3000=9000
How much did Kevin invest in Fund A? $9,000